OpenStudy (anonymous):

The value of a two-digit number is twice as large as the sum of its digits. If the digits are reversed, the resulting number would be 9 less than 5 times the original number. Find the original number.

OpenStudy (anonymous):

The value of a two digit number is twice as large as the sum of its digits. If the digits were reversed, the resulting number would be 9 less than 5 times the original number. Find the original number. t = tens digit u = units digit 10t+u = original number 10u+t = the number named when the digits were reversed 10u+t = 9-5(10t+u) ------------------ EQUATION: 10t + u = 2(t + u) 10u + t = 5(10t+u)-9 ------------------------- Rearrange the equations: 8t -u = 0 49t - 5u = 9 ------------------ solve by substitution: 49t - 5(8t) = 9 9t = 9 t = 1 ---------- Substitute to solve for "u": u = 8t so u = 8 ------------------- Original Number 18 =======================

OpenStudy (anonymous):

you are amazing and u do it exactly the way I do u get a medal and are a fav!!!

OpenStudy (anonymous):

Thanks @Boyce Glad I could help! :-)

OpenStudy (anonymous):

How old are u? I am almost 16.